OpenStudy (anonymous):

Does the equation y= 1+ radical x have a minimum value or a maximum value?

OpenStudy (anonymous):

x=0 y=1+0^(1/2)=1 - min

OpenStudy (anonymous):

we know\[\sqrt{x}\ge 0\]so\[y=1+\sqrt{x}\ge1\]